# Mingkai Liao

### v1

@@ 4.2 Two-Echo-Chamber Training Instance Distribution
PACIFIER follows an inductive learning paradigm and is trained offline on a distribution of synthetic graph–opinion instances. The purpose of this distribution is not to exactly replicate any single real-world network, but to expose the agent to a diverse yet structured set of polarized states that capture the dominant regimes observed in practice.
@@
Across datasets, we observe that these statistics consistently fall within the range spanned by the synthetic BA-based two-group graphs generated by $\mathcal{D}$. In particular, real graphs exhibit (i) strong within-camp connectivity, (ii) sparse but non-negligible cross-camp exposure, and (iii) degree heterogeneity compatible with preferential-attachment structure. These observations indicate that the proposed training distribution captures the dominant polarization regimes present in real data, thereby narrowing the train–test distribution gap and improving generalization.
@@ 4.3 Encoding
Definition 4.1 (Topology-preserving intervention).
An intervention process is topology-preserving if, throughout the decision horizon, the graph topology remains exactly unchanged: no nodes or edges are removed, added, rewired, or reweighted, and hence no node or edge becomes structurally inactive. Formally, the graph structure $G=(V,E)$ is invariant (i.e., $V$ and $E$ do not change with $t$), and the intervention only modifies node/edge attributes or imposes constraints (e.g., internal opinions, expressed opinions, or fixed-value boundary conditions). In such settings, the adjacency structure alone cannot reveal which nodes have been intervened on, creating a representation ambiguity under permutation-invariant message-passing GNN encoders, where different intervention histories can produce nearly identical aggregated embeddings, leading to state aliasing and degraded value estimation in reinforcement learning.
This topology-preserving setting gives rise to two fundamental representation challenges: (i) history-induced state aliasing under a fixed topology, and (ii) the lack of polarization-related global signals in standard message-passing representations, which can hinder accurate value estimation and stable policy learning when polarization is driven by nontrivial interactions between opinion states and network structure. PACIFIER addresses these challenges via two encoding innovations: a temporal-aware node marking mechanism to explicitly encode intervention history, and polarization-related auxiliary features that summarize opinion–structure interaction patterns to provide value-relevant global cues.
State.
At decision step $t$, the environment state is represented as
$s_{t}=(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t}),$
where $G=(V,E)$ is the (fixed) network, $X_{t}\in\mathbb{R}^{n\times d_{0}}$ is the node-feature matrix, $\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$ is a graph-level auxiliary feature vector, and $\mathbf{m}_{t}\in\{0,1\}^{n}$ is a feasibility mask indicating which nodes can still be selected for intervention.
Encoding pipeline overview (graph + node features + super-node aggregation).
Figuratively, PACIFIER’s encoder takes as input (i) the fixed adjacency structure of $G$ and (ii) per-node attributes that evolve with interventions, and produces node embeddings and a graph-level embedding that are reused to score all actions in parallel. The pipeline is summarized as follows.
$\mathbf{h}_{v}^{(0)}=\sigma\!\left(W_{0}\mathbf{x}_{t}(v)\right),\qquad\mathbf{h}_{v}^{(0)}\leftarrow\mathrm{norm}\!\left(\mathbf{h}_{v}^{(0)}\right),$ (8)
where $\mathrm{norm}(\cdot)$ denotes $\ell_{2}$ normalization. This corresponds to the input projection layer in our implementation (the $w_{n2l}$ mapping).
$\mathbf{a}_{v}^{(k)}$ $=\sum_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(k-1)},$ (9)
$\mathbf{h}_{v}^{(k)}$ $=\sigma\!\left(W_{\mathrm{sage}}\begin{bmatrix}W_{\mathrm{nbr}}\mathbf{a}_{v}^{(k)}\\ W_{\mathrm{self}}\mathbf{h}_{v}^{(k-1)}\end{bmatrix}\right),\qquad\mathbf{h}_{v}^{(k)}\leftarrow\mathrm{norm}\!\left(\mathbf{h}_{v}^{(k)}\right).$ (10)
This design matches the implementation pattern: a sparse neighbor-sum (n2nsum_param) followed by linear transforms (p_node_conv, p_node_conv2) and a concatenation-based update (p_node_conv3).
$\mathbf{g}^{(k)}=\sum_{v\in V}\mathbf{h}_{v}^{(k)},$ (11)
which provides an iteration-synchronized global summary of the evolving opinion/intervention state. In our implementation, this is realized by an explicit aggregation operator (a sparse subgsum_param that pools node embeddings into a batch-level vector), conceptually equivalent to a virtual node connected to all nodes and updated by one-way aggregation.
Node features (temporal-aware marking; Innovation 1).
Let $S_{t}\subseteq V$ denote the set of nodes that have already been intervened by step $t$. In topology-preserving moderation, the adjacency structure alone does not encode intervention history; consequently, different intervention sequences may produce graph states that are indistinguishable if history is not explicitly represented. This state aliasing can severely degrade value estimation and credit assignment in sequential decision making.
To resolve this issue, PACIFIER explicitly encodes intervention history into node features via a temporal-aware marking mechanism. Each node $v$ is assigned a feature vector
@@
Polarization-related auxiliary features (Innovation 2).
While learned embeddings capture rich local and global structure, polarization moderation objectives depend explicitly on how the opinion configuration interacts with the network. Purely message-passing representations can be insufficient to expose the global signals that correlate with polarization regimes (e.g., cross-camp exposure among remaining active nodes, or the level of within-camp cohesion), especially under one-shot planning where the policy input should not rely on repeated steady-state recomputation.
We therefore augment learned embeddings with a deterministic auxiliary feature vector $\mathbf{u}_{t}$ that summarizes intervention progress and opinion–structure interaction patterns. Concretely, let $n=|V|$ and $m=|E|$. We include the covered-node ratio
@@
Output of encoding.
The encoding stage produces (i) node embeddings $\{\mathbf{h}_{v}\}_{v\in V}$, (ii) a graph-level embedding $\mathbf{g}_{t}$ (via super-node/global aggregation), (iii) polarization-related auxiliary features $\mathbf{u}_{t}$, and (iv) a feasibility mask $\mathbf{m}_{t}$. These representations are then passed to the decoding module to score all candidate intervention actions.
@@ 4.5 Flexibility
  • FJ-model variants, including non-analytic nonlinear dynamics (e.g., biased assimilation). PACIFIER is agnostic to the specific analytical form of the underlying opinion dynamics. Any FJ-consistent variant—including nonlinear extensions such as biased assimilation that typically admit no closed-form steady state—can be incorporated by modifying only the environment transition and polarization evaluation, while keeping the same encoder–decoder architecture and learning procedure unchanged.
Overall, this modularity allows PACIFIER to serve as a unified learning-based intervention framework for a broad class of polarization moderation problems, spanning different opinion representations, intervention mechanisms, and dynamical models within the Friedkin–Johnsen paradigm and beyond.
@@ 5 Experiments
In this section, we evaluate the proposed method on real-world social network datasets with explicit opposing camps. Our experiments are designed to answer the following questions: (i) whether the proposed approach can effectively moderate polarization in realistic network settings, and (ii) how its performance compares to baseline strategies under the same intervention budget. We begin by describing the datasets used in our evaluation, followed by the experimental setup, baselines, and evaluation metrics.
@@ 5.1 Datasets
Figure 1: Dataset-level relationships between initial polarization $\pi(\mathbf{z}^{(0)})$ and two structural indicators across the 31 topic graphs. Higher cross-camp mixing (larger $\rho_{\pm}$) and higher connectivity (larger $\bar{d}$) are associated with lower initial polarization. Pearson and Spearman statistics are reported in the plot titles.
@@
While Table 1 summarizes all 31 structurally valid topic graphs, our main experiments further focus on instances with sufficiently strong initial polarization. This additional filtering is motivated by two considerations. First, when the initial polarization is very small, the achievable reduction under any intervention budget is limited, which yields weak reward signals and makes reinforcement-learning-based methods difficult to train and evaluate fairly. Second, we aim to select a threshold that retains the majority of genuinely polarized scenarios while excluding graphs that are clearly weakly polarized. Consistent with the dataset-level correlation analysis in Fig. 1(a)1(c), higher initial polarization tends to coincide with more pronounced echo-chamber structure (i.e., lower cross-camp exposure $\rho_{\pm}$ and typically lower overall connectivity), so focusing on sufficiently polarized instances also emphasizes regimes where echo chambers are structurally clearer.
@@
In addition to the real-world benchmark, we also summarize the structural regimes covered by our synthetic two-echo-chamber graphs used for offline training and evaluation. Following Sec. 4.2, we generate BA-based two-community graphs with sparse inter-community edges and assign extreme camp opinions ($\pm 1$) according to the planted partition. We report, for each size range, the number of generated graphs and the observed value ranges (min–max across graphs) of three indicators: average degree $\bar{d}$, cross-camp edge ratio $\rho_{\pm}$, and the resulting initial polarization $\pi(\mathbf{z}^{(0)})$.
As summarized in Table 4, the synthetic graphs span a broad yet consistent regime across all size ranges. Connectivity ranges from sparse to moderately dense graphs (approximately $\bar{d}\in[1.87,11.66]$ across all ranges), while cross-camp exposure remains low by construction (typically $\rho_{\pm}\in[0.00,0.17]$), producing strongly polarized initial states with $\pi(\mathbf{z}^{(0)})$ often close to $1.0$ and ranging roughly from $0.06$ to $1.00$ depending on the sampled inter-camp edges. Importantly, these ranges cover the regimes emphasized by our real-world filtered benchmark (Table 3), where echo-chamber structure corresponds to low $\rho_{\pm}$ and higher polarization. Moreover, the per-range statistics are highly stable as graph size increases from 30–50 up to 400–500 nodes: the min–max intervals of $\bar{d}$, $\rho_{\pm}$, and $\pi(\mathbf{z}^{(0)})$ remain comparable across ranges. This indicates that these three indicators are largely scale-insensitive in our synthetic generator—they characterize the echo-chamber regime (connectivity and mixing) rather than the absolute graph size—which is desirable for inductive learning and evaluation across sizes. In particular, the training range (30–50) already overlaps the global synthetic ranges and exhibits the same low-mixing/high-polarization pattern, while larger ranges preserve similar statistics, supporting inductive generalization from small training graphs to larger test graphs.
Table 4: Per-range statistics of synthetic graphs used for training/testing. For each node-size range, we report the number of generated graphs and the min–max range of $\bar{d}$, $\rho_{\pm}$, and $\pi(\mathbf{z}^{(0)})$ under the extreme two-camp assignment.

### v2

@@ 4.1 Two-Echo-Chamber Training Instance Distribution
PACIFIER follows an inductive learning paradigm and is trained offline on a distribution of synthetic graph–opinion instances. The purpose of this distribution is not to exactly replicate any single real-world network, but to expose the agent to a diverse yet structured set of polarized states that capture dominant regimes observed in practice.
@@
Across datasets, we observe that these statistics consistently fall within the range spanned by the synthetic BA-based two-group graphs generated by $\mathcal{D}$. In particular, real graphs exhibit (i) strong within-camp connectivity, (ii) sparse but non-negligible cross-camp exposure, and (iii) degree heterogeneity compatible with preferential-attachment structure. These observations indicate that the proposed training distribution captures dominant polarization regimes present in real data, thereby narrowing the train–test distribution gap and improving generalization.
@@ 4.2 Encoding
(Innovation 1) Node Features (Temporal-aware marking).
Let $S_{t}\subseteq V$ denote the set of nodes that have already been intervened by step $t$. In topology-preserving moderation, the adjacency structure alone does not encode intervention history; consequently, different intervention sequences may produce graph states that are indistinguishable if history is not explicitly represented. To resolve this issue, PACIFIER explicitly encodes intervention history into node features via a temporal-aware marking mechanism. Each node $v$ is assigned a feature vector
@@
(Innovation 2) Polarization-related Auxiliary Features.
PACIFIER augments learned embeddings with a deterministic auxiliary feature vector $\mathbf{u}_{t}$ that summarizes intervention progress and opinion–structure interaction patterns. Concretely, let $n=|V|$ and $m=|E|$. We include the covered-node ratio
@@
Encoding Pipeline Overview.
PACIFIER’s encoder takes as input (i) the fixed adjacency structure of $G$ and (ii) per-node attributes that evolve with interventions, and produces node embeddings and a graph-level embedding that are reused to score all actions in parallel.
$\mathbf{h}_{v}^{(0)}=\sigma(W_{0}\mathbf{x}_{t}(v)),\qquad\mathbf{h}_{v}^{(0)}\leftarrow\mathrm{norm}(\mathbf{h}_{v}^{(0)}),$ (14)
where $\mathrm{norm}(\cdot)$ denotes $\ell_{2}$ normalization.
$\mathbf{a}_{v}^{(k)}$ $=\sum_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(k-1)},$ (15)
$\mathbf{h}_{v}^{(k)}$ $=\sigma\left(W_{\mathrm{sage}}\begin{bmatrix}W_{\mathrm{nbr}}\mathbf{a}_{v}^{(k)}\\ W_{\mathrm{self}}\mathbf{h}_{v}^{(k-1)}\end{bmatrix}\right),\qquad\mathbf{h}_{v}^{(k)}\leftarrow\mathrm{norm}(\mathbf{h}_{v}^{(k)}).$ (16)
$\mathbf{g}^{(k)}=\sum_{v\in V}\mathbf{h}_{v}^{(k)},$ (17)
which provides an iteration-synchronized global summary of the evolving opinion/intervention state.
@@ 4.3 Decoding
4.4 Method Procedure
We now describe the end-to-end decision-making procedure of PACIFIER, including the MDP formulation, offline training of the two variants, and the deployment-time feed-forward execution on a target network. Crucially, the policy input is always the compact representation $(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t})$ produced by the encoder, so deployment-time planning does not require iterative steady-state recomputation.
MDP formulation.
State.
At step $t$, the state is
$s_{t}=(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t}),$
where $G=(V,E)$ is the (fixed) network, $X_{t}\in\mathbb{R}^{n\times d_{x}}$ is the node-feature matrix, $\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$ is a graph-level auxiliary feature vector, and $\mathbf{m}_{t}\in\{0,1\}^{n}$ is an action-feasibility mask. Let $S_{t}\subseteq V$ denote the set of already intervened nodes. Then $\mathbf{m}_{t}(v)=0$ for $v\in S_{t}$ and $\mathbf{m}_{t}(v)=1$ otherwise. Under MI/ME-style moderation, the topology of $G$ remains fixed, and state evolution is driven by opinion updates and intervention history encoded in $X_{t}$ (Sec. 4.2).
Action.
An action selects one feasible node:
$a_{t}\in\mathcal{A}_{t}=\{v\in V\mid\mathbf{m}_{t}(v)=1\}.$
Transition.
Reward (trajectory-aware signal).
$r_{t}=-\,\frac{\pi(\mathbf{z}^{(t)})}{C}\cdot c(a_{t}).$
For unweighted settings, $c(v)\equiv 1$.
Policy parameterization via encode–decode.
Offline training.
PACIFIER-RL. During offline training, PACIFIER-RL interacts with sampled synthetic instances from $\mathcal{D}$ (Sec. 4.1) to generate trajectories $\{(s_{t},a_{t},r_{t},s_{t+1})\}$ over a fixed horizon (budget) $k$ (or until no feasible action remains). At each step, the agent observes the state representation $(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t})$ and selects one feasible node via an $\epsilon$-greedy strategy: with probability $1-\epsilon$ it chooses the feasible node with the largest predicted $Q_{\theta}(s_{t},v)$, and otherwise samples a feasible node uniformly at random. The exploration rate $\epsilon$ is annealed from a high initial value to a small final value over training episodes.
$R_{t:t+n}=\sum_{i=0}^{n-1}\gamma^{i}r_{t+i}.$
$y_{t}=R_{t:t+n}+\gamma^{n}\max_{v}Q_{\theta^{-}}(s_{t+n},v),$
$y_{t}=r_{t}.$
Parameters are optimized by minimizing
Online application (deployment).
After offline training, both variants are deployed on a target real-world graph in a purely feed-forward manner. At each decision step $t$, we construct the current state representation $(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t})$ (Sec. 4.2) and evaluate $Q_{\theta}(s_{t},v)$ for all nodes in parallel (Sec. 4.3). Feasibility is enforced by the mask $\mathbf{m}_{t}$: nodes with $\mathbf{m}_{t}(v)=0$ (already intervened) are assigned a large negative value and are never selected. The deployed policy is greedy with respect to $Q_{\theta}$:
@@ 4.5 Flexibility
Overall, this modularity allows PACIFIER to serve as a unified learning-based intervention framework for a broad class of polarization moderation problems within the Friedkin–Johnsen paradigm, covering binary and continuous opinions, unweighted and cost-weighted interventions, as well as both topology-preserving and topology-altering settings.
@@ 5 Experiments
In this section, we conduct a comprehensive empirical evaluation of PACIFIER on both synthetic and real-world polarized social networks. Our experiments are designed to answer four questions: (i) how PACIFIER performs against classical baselines on the canonical moderation tasks, (ii) whether the learned policies generalize across graph sizes and from synthetic training graphs to heterogeneous real-world networks, (iii) how PACIFIER behaves in topology-preserving versus topology-altering intervention settings, and (iv) whether the proposed temporal-aware marking and polarization-aware auxiliary features contribute to performance.
To answer these questions, we first describe the datasets, filtering criteria, evaluation protocol, and baselines. We then report results on synthetic benchmarks and real-world datasets, followed by runtime comparisons and an ablation study.
@@ 5.1 Datasets
Figure 1: Dataset-level relationships between initial polarization $\pi(\mathbf{z}^{(0)})$ and two structural indicators across the 31 topic graphs. Higher cross-camp mixing (larger $\rho_{\pm}$) and higher connectivity (larger $\bar{d}$) are associated with lower initial polarization. Pearson and Spearman statistics are reported in the plots.
@@
While Table 1 summarizes all 31 structurally valid topic graphs, our main experiments further focus on instances with sufficiently strong initial polarization. This additional filtering is motivated by two considerations. First, when the initial polarization is very small, the achievable reduction under any intervention budget is limited, which yields weak reward signals and makes learning-based methods difficult to train and evaluate fairly. Second, we aim to select a threshold that retains the majority of genuinely polarized scenarios while excluding graphs that are clearly weakly polarized. Consistent with the dataset-level correlation analysis in Fig. 1(a)1(c), higher initial polarization tends to coincide with more pronounced echo-chamber structure (i.e., lower cross-camp exposure $\rho_{\pm}$ and typically lower overall connectivity), so focusing on sufficiently polarized instances also emphasizes regimes where echo chambers are structurally clearer.
@@
In addition to the real-world benchmark, we also summarize the structural regimes covered by our synthetic two-echo-chamber graphs used for offline training and evaluation. Following Sec. 4.1, we generate two-community BA graphs, assign extreme camp opinions ($\pm 1$) according to the planted partition, and then add sparse inter-community edges. Compared with the earlier synthetic setting, the current generator further applies an initial-polarization filtering step: a sampled graph is retained only if its initial polarization satisfies $\pi(\mathbf{z}^{(0)})\geq 0.4$. This ensures that all synthetic instances used in training and testing lie in a sufficiently polarized regime and avoids weakly polarized cases with limited room for meaningful intervention.
For each node-size range, we report the number of generated graphs and the observed value ranges (min–max across graphs) of three indicators: average degree $\bar{d}$, cross-camp edge ratio $\rho_{\pm}$, and the resulting initial polarization $\pi(\mathbf{z}^{(0)})$.
As summarized in Table 4, the filtered synthetic graphs remain structurally consistent across all size ranges. The average degree now lies in a moderately denser regime, roughly from $3.56$ to $15.81$, while cross-camp exposure remains low by construction, typically between $0.00$ and $0.12$. Due to the explicit rejection step enforcing $\pi(\mathbf{z}^{(0)})\geq 0.4$, all retained graphs exhibit clearly polarized initial states, with observed polarization ranges concentrated in $[0.40,1.00]$ up to minor range-specific variation.
Importantly, these filtered synthetic regimes are more closely aligned with the real-world filtered benchmark used in our main experiments, where we also focus on sufficiently polarized instances. Moreover, the per-range statistics remain highly stable as graph size increases from 30–50 to 400–500 nodes: although the upper bound of $\bar{d}$ grows moderately with size, the ranges of $\rho_{\pm}$ and $\pi(\mathbf{z}^{(0)})$ stay remarkably consistent. This suggests that the synthetic generator mainly controls the polarization regime—namely, within-camp connectivity, limited cross-camp mixing, and sufficiently high initial polarization—rather than producing size-specific distribution shifts. In particular, the training range (30–50) already exhibits the same low-mixing and high-polarization pattern as the larger test ranges, supporting inductive generalization from small training graphs to larger synthetic graphs.
Table 4: Per-range statistics of synthetic graphs used for training/testing. For each node-size range, we report the number of generated graphs and the min–max range of $\bar{d}$, $\rho_{\pm}$, and $\pi(\mathbf{z}^{(0)})$ under the extreme two-camp assignment after applying the initial-polarization filter $\pi(\mathbf{z}^{(0)})\geq 0.4$.
@@ 5.2 Visualization Protocol and Evaluation Views
5.3 Baseline Methods
We compare PACIFIER against a set of analytical, heuristic, structural, and oracle-style baselines. Unless otherwise stated, all methods without the suffix -FI operate under the standard one-shot planning regime defined in Sec. 3.4: they must commit to a complete intervention sequence using only the initial instance information and are not allowed to use recomputed intermediate settled-opinion states during planning. Methods with the suffix -FI are their full-information counterparts: they are allowed to access recomputed settled-opinion states after each partial intervention step and can therefore replan sequentially using intermediate feedback.
PACIFIER variants.
We evaluate two versions of the proposed method. PACIFIER-RL is the full reinforcement-learning variant, which learns long-horizon action values and performs sequential intervention by greedy decoding of the learned $Q$-function. PACIFIER-Greedy shares the same encoder–decoder architecture but is trained only on immediate rewards, thus serving as a myopic learned ranking policy. Both variants follow the one-shot regime in the standard experiments. In the full-information comparison, we additionally consider PACIFIER-RL-FI and PACIFIER-Greedy-FI, which expose recomputed intermediate settled-opinion information to the same learned policies during sequential selection.
BOMP.
BOMP (Binary Orthogonal Matching Pursuit) is the main analytical baseline for MI, following Matakos et al. [8]. It exploits the linear structure of the FJ steady-state solution and the associated global influence operator to iteratively select nodes whose internal-opinion neutralization is expected to most reduce polarization. BOMP is a strong reference in strictly linear MI settings, but its dense global computations become costly on large graphs and it is not naturally suited to topology-altering settings such as node removal.
ExtremeExpressed.
ExtremeExpressed is a lightweight opinion-based heuristic that ranks candidate nodes according to the extremeness of their current expressed opinions. Intuitively, it prioritizes nodes whose expressed states are most polarized in magnitude, on the assumption that directly moderating such nodes yields strong immediate effect. In the standard one-shot setting, the ranking is determined from the initial settled-opinion state only. Its full-information counterpart, ExtremeExpressed-FI, recomputes the settled-opinion state after each intervention step and re-ranks the remaining candidates accordingly.
ExtremeNeighbours.
ExtremeNeighbours is another opinion-aware heuristic that scores a node using the extremeness of its local neighborhood, aiming to identify nodes embedded in strongly polarized local environments. Compared with ExtremeExpressed, it injects a limited amount of neighborhood structure into the ranking rule, but still remains a simple non-learning baseline. As above, the standard version is one-shot, whereas ExtremeNeighbours-FI is the full-information version that recomputes the intermediate settled state and updates the ranking after each intervention step.
PageRank.
PageRank is a purely structural baseline that ranks nodes by global graph centrality, independent of the current opinion state. It provides a strong topology-only reference for testing whether structural importance alone is sufficient for effective moderation. Because it does not depend on intermediate settled-opinion recomputation, we use PageRank only in its standard one-shot form in the main comparisons.
Random.
Random selects nodes uniformly at random from the feasible set. It serves as a minimal baseline and helps quantify the extent to which all other methods improve over uninformed intervention.
Cost.
Cost is a simple cost-only baseline used exclusively in the cost-weighted intervention tasks. It ranks all feasible nodes in ascending order of their intervention cost and directly selects the cheapest nodes as the intervention sequence, without using any opinion or structural information. This baseline serves to test whether performance in cost-aware settings can be explained merely by preferring low-cost nodes, rather than by reasoning jointly about cost, network structure, and opinion dynamics. Because its ranking depends only on the static node costs, we use Cost only in the standard one-shot form.
Greedy (oracle-style full-information baseline).
In the synthetic full-information comparison, we additionally include Greedy, an oracle-style replanning baseline. At each step, it exhaustively tries every feasible candidate node, temporarily applies the intervention, recomputes the resulting settled-opinion state, evaluates the immediate polarization reduction, rolls the intervention back, and then selects the best candidate. This method explicitly uses intermediate settled-opinion information and therefore belongs to the full-information regime. It is computationally expensive but provides a strong upper-level myopic reference for understanding the value of sequential replanning with explicit feedback.
One-shot vs. full-information distinction.
The distinction between one-shot and full-information baselines is central to our evaluation. One-shot methods are constrained to plan from the initial instance only and therefore match the practical deployment regime targeted by PACIFIER. By contrast, full-information methods are allowed to repeatedly recompute intermediate settled-opinion states and update their decisions online. Our standard experiments compare PACIFIER against one-shot baselines, while the dedicated full-information synthetic benchmark separately examines how much additional benefit can be obtained when such intermediate feedback is made available.